18.099b Problem Set 1 Due: Thursday, February 18th (in class or before).

1. Suppose x is a real number greater than 1. Show that for every real number y there exists a positive integer n such that x^{n} > y. (Hint. Use the least upper bound property of the real numbers: Every nonempty bounded set of real numbers has a least upper bound.)

2. Latex exercise.